Can you please help?
(a^-1/3b^1/2)^4(a^-1/2b^3/5)^10
I'm lost!

Answer by scott8148(5880)

exponents raised to an exponent are multiplied ... (a^b)^c=a^(bc)

when equal bases are multiplied, the exponents are added ... (a^b)(a^c)=a^(b+c)

(a^(-4/3)b^2)(a^(-5)b^6)=a^(-19/3)b^8